# Physics-informed deep learning for infectious disease forecasting

### v1

@@ 5.1 Impact of training data length
We trained PINNs using data from the most recent 4, 8, 12, 16, 20, and 24 weeks and evaluated their predicting performances on weeks 48 to 110. The result is shown in Fig. 5. Each subplot shows the MASE scores of predictions on cases, deaths, and hospitalizations for the following 1, 2, 3, and 4 weeks by using different lengths of training data ranging from 4 to 24 weeks. For all the three targets, the overall trend of MASE values is that as the training data length becomes longer, MASE gradually falls and then becomes stable. When we only use recent data from 4 or 8 weeks, it seems the data is not enough to capture the disease dynamics. When we use data from the past 12 or 16 weeks, it could present a comparable accuracy compared with PINNs using all the historical data. We also note that when the training data keeps increasing to the recent 20 or 24 weeks’ data, the accuracy of the performance doesn’t make significant further improvement, which is likely that these training data span nearly half a year, and thus they are not closely related to current epidemiology trends and don’t have a significant contribution to predictions.
@@ 5.2 Impact of L2 regularization
We further show that without L2 regularization, the model accuracy drops significantly. Removing the L2 regularization results in average decreases of 22.48% in cases, 9.16% in deaths, and 42.54% in hospitalizations for the following 1-4 weeks. This result suggests that L2 regularization can still be beneficial to model predictions in the PINNs model.

### v2

@@ 5.1 Impact of training data length
We trained PINNs using datasets comprising the most recent 4, 8, 12, 16, 20, and 24 weeks of observations, and evaluated their forecasting performance over the period spanning weeks 48 to 110. The results, shown in Fig. 7, illustrate the MASE scores for predictions of cases, deaths, and hospitalizations across 1-, 2-, 3-, and 4-week horizons using different training data lengths.
For all three targets, we observe a consistent trend: as the length of training data increases, the MASE scores generally decrease and then stabilize. Models trained on only 4 or 8 weeks of data tend to underperform, likely due to insufficient information to accurately capture the underlying disease dynamics. In contrast, models trained on 12 or 16 weeks of data achieve predictive accuracy comparable to those trained on the full historical dataset. Notably, further increasing the training window to 20 or 24 weeks yields only marginal improvements in accuracy. This suggests that data beyond 16 weeks—spanning nearly half a year—may be less relevant to current epidemiological conditions and contribute less to near-term forecasting performance.
@@ 5.2 Impact of L2 regularization
We further demonstrate that the absence of L2 regularization leads to a substantial decline in model accuracy. Specifically, removing L2 regularization results in average performance drops of 22.48% for cases, 9.16% for deaths, and 42.54% for hospitalizations over the 1- to 4-week forecasting horizons. These results highlight the importance of incorporating L2 regularization in the PINNs framework, as it contributes to improved generalization and prediction stability across multiple epidemiological targets.